There is a rotor ride at Hershey Park where the radius of the ride is 10 meters. If the coefficient of friction of the walls of the ride is 2.5, what is the velocity needed for the rotator ride for you not to fall down the wall.


1
Expert's answer
2021-03-21T11:23:23-0400
"F_c=F_{fr},""\dfrac{mv^2}{R}=\mu mg,""v=\sqrt{\mu gR}=\sqrt{2.5\cdot9.8\ \dfrac{m}{s^2}\cdot10\ m}=15.65\ \dfrac{m}{s}."